The perimeter of a rectangle is 70 cm. The length of the rectangle is 4 cm less than double the width. Find the length and width of the rectangle
Answer by mananth(16949)

let width =x
length = 2x-4
perimeter = 70
..
2*(L+W)=perimeter
2*(x+2x-4)=70
/2
3x-4=35
+4
3x=39
/3
x=13 the width
length = 2x-4===>2*13-4=22
m.ananth@hotmail.ca